OpenStudy (anonymous):

Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0

OpenStudy (anonymous):

Solution: 10 only solution

OpenStudy (saifoo.khan):

Look, any problem like this

OpenStudy (saifoo.khan):

x = 10 & x = 10

OpenStudy (saifoo.khan):

x-10 = 0 & x -10 = 0 (x-10)(x-10) = 0 FOIL.